OpenStudy (anonymous):

Between what two consecutive integers does the square root of 155 lie?

OpenStudy (anonymous):

12*12 = 144 I bet that gets close

OpenStudy (anonymous):

i would just take the sqrt on a calculator and see where it falls. right?

OpenStudy (unklerhaukus):

\[12^2=144\] \[13^2=169\]

OpenStudy (anonymous):

144 and 169 are the two #'s it falls between?

OpenStudy (anonymous):

no because they are not consecutive, meaning on after the other like 1 and 2, or 55 and 56

OpenStudy (anonymous):

These are the answers A.11 and 12 B.12 and 13 C. 13 and 14 D.14 and 15

OpenStudy (anonymous):

it would be the 2 number that @UnkleRhaukus pointed out

OpenStudy (anonymous):

K thanks :)

OpenStudy (anonymous):

youre welcome

OpenStudy (unklerhaukus):

\[\sqrt{144}<\sqrt {155}<\sqrt169\]